OpenStudy (anonymous):

factorize (2^n+1 + 2^n )/3, hence write : (2^1001+2^1000)/3 as a power of 2

OpenStudy (mertsj):

\[2^{n+1}+2^{n}=2^{n}(2^{1}+1)=2^{n}(3)\]

OpenStudy (anonymous):

=(2^n(2^1)+2^n)/3 =(2^n(3))/3 =3(2^n)/3 =2^n Therefore, when n=1000 2^n= 2^1000= 1.07150861 × 10^301

OpenStudy (mertsj):

And \[\frac{2^{n}(3)}{3}=2^{n}\]

OpenStudy (mertsj):

So \[\frac{2^{1001}+2^{1000}}{3}=2^{1000}\]

OpenStudy (anonymous):

ok, i am understand now....

OpenStudy (mertsj):

Good.